OpenStudy (anonymous):

What is the solution to the system of equations? x+y=5 x-2y=2 (___,___)

OpenStudy (anonymous):

Please Help!!

OpenStudy (solomonzelman):

You can rearrange the second equation: \(\large\color{black}{ \displaystyle x-2y=2 }\) \(\large\color{black}{ \displaystyle x-2y\color{blue}{+2y}=2\color{blue}{+2y} }\) \(\large\color{black}{ \displaystyle x=2+2y }\) And the substitute this for x into your first equation: \(\large\color{black}{ \displaystyle \color{red}{x}+y=5 }\) \(\large\color{black}{ \displaystyle \color{red}{\left(2+2y\right)}+y=5 }\)

OpenStudy (solomonzelman):

Can you solve for y from here?

OpenStudy (anonymous):

umm, I'm like clueless when It comes to this stuff lol I'm sorry

OpenStudy (solomonzelman):

Or perhaps you can find some other sources on solving a system of linear equations, by the method of: 1) «Substitution» 2) «Elimination»

OpenStudy (solomonzelman):

good Luck

OpenStudy (anonymous):

Thanks